OpenStudy (freckles):

@Loser66 fun question Parametrize the line segment \[y=2x \text{ where } x \in [1,3] \text{ on the interval } t \in [a,b]\]

OpenStudy (freckles):

I made up this question. Hopefully the wording is correct. I think it is.

OpenStudy (astrophysics):

Could we just use x=a+bt, y=d+te, where we start at x = a and y = d, then move in the direction of (b,e) from what I remember at least when parametrize lines xD

OpenStudy (fibonaccichick666):

x=t then y=2t :)

OpenStudy (astrophysics):

Yeah, x = t, then y = g(t)

OpenStudy (fibonaccichick666):

hehe, yep. the easy way out

OpenStudy (freckles):

This is what I did: \[x(t)=ct+d \\ 1 \le ct+d \le 3 \\ \text{ assume } c>0 \\ \frac{1-d}{c} \le t \le \frac{3-d}{c} \\ \text{ choose } a=\frac{1-d}{c} \\ \text{ and } b=\frac{3-d}{c}\] \[ac=1-d \\ bc=3-d \\ \text{ subtract these equations giving } \\ (a-b)c=-2 \\ c=\frac{-2}{a-b}=\frac{2}{b-a} , \text{ then go back and solve for } d \text{ in terms of } a \text{ and } b \] \[ac=1-d \\ ac-1=-d \\ d=1-ac \\ d=1-a \cdot \frac{2}{b-a}=1-\frac{2a}{b-a} \\ \text{ so we have } \\ x(t)=\frac{ 2 t}{b-a}+(1-\frac{2a}{b-a}) \\ \text{ or } \\ x(t)=\frac{2}{b-a}(t-a)+1 \\ \text{ and so } \\ y(t)=\frac{4}{b-a}(t-a)+2 \\ (x,y)=(\frac{2}{b-a}(t-a)+1, \frac{4}{b-a}(t-a)+2) \text{ where } t \in [a,b]\]

OpenStudy (freckles):

I think x(t)=t and y(t)=2t would have worked if I wanted t to be between 1 and 3 like x is

OpenStudy (fibonaccichick666):

oh, you want the limits as t goes to a to equal the corresponding things?

OpenStudy (freckles):

yes I wanted t to be between a and b

OpenStudy (fibonaccichick666):

ah, I thought we got to pick a and b.

OpenStudy (astrophysics):

That is very cool!

OpenStudy (freckles):

It is just kind of fun. And I just wanted to see if @Loser66 could get there. :p

OpenStudy (astrophysics):

Ah yeah, sorry about intruding

OpenStudy (freckles):

It is fine. He isn't here. I'm glad you guys came so I could have someone to play with.

OpenStudy (astrophysics):

I'm looking at that pdf, the archimedes spiral looks very fun

OpenStudy (freckles):

Looks fun, yep. We didn't go that far. I think I seen that spiral thing before somewhere... Not sure I thought I remember something about Fibonacci sequence.

OpenStudy (fibonaccichick666):

yea, if illustrating the fibonacci sequence you can use a spiral

OpenStudy (astrophysics):

Reminds me of @FibonacciChick666

OpenStudy (astrophysics):

hehe yes

OpenStudy (fibonaccichick666):

here check it out, https://en.wikipedia.org/wiki/Golden_spiral

OpenStudy (fibonaccichick666):

:D

OpenStudy (freckles):

Look at you @FibonacciChick666 you are famous

OpenStudy (freckles):

I guess those spirals are a little different.

OpenStudy (fibonaccichick666):

what can I say, I chose my name well. and yea, that's an archimedes spiral

OpenStudy (freckles):

when I google freckles I think I see a lot more results come up then when I google fibonacci though :p

OpenStudy (freckles):

freckles=19,700,000 fibonacci=16,500,000 freckles>fibonacci

OpenStudy (freckles):

you beat astrophysics though astrophysics=14,400,000

OpenStudy (astrophysics):

Haha :c

OpenStudy (freckles):

dan815 comes in last place dan815=20,900

OpenStudy (fibonaccichick666):

lol

OpenStudy (baru):



OpenStudy (baru):

this is how i did it :) we travel from A to B in (b-a) seconds, so unit vector with w.r.t time = AB/(b-a) =< 2, 4> /(b-a) Now we parametrize with t' and velocity (b-a) parametirc eq C(t' )=\[C=~~<1,2> + \frac{ <2,~4 > }{ b-a }t'\] but t'=t-a because 'a' seconds have already passed when we reach pt A. (same logic as shifting origin in graphs thus \[C=~~<1,2> + \frac{ <2,~4 > }{ b-a }(t-a)\]

OpenStudy (baru):

*<2,4> is vector AB ( i forgot to mention)

OpenStudy (freckles):

cute stuff

OpenStudy (baru):

:p

OpenStudy (freckles):

it might seem easier than what I did :p

OpenStudy (freckles):

jk it is

OpenStudy (baru):

* when i said "unit vector w.r.t time" i meant "velocity vector" :( i worded this very poorly, sorry

OpenStudy (freckles):

its k I think I'm a bad worder too